Please answer with solution thank you so much. Answer with solution. Write the balance equation of the reaction when 1 gram of iron filings was to be reacted with 20 ML 3.0 M hydrochloric acid and the gas collected in a container. The limiting reactant is? What gas is formed, calculate its no. of moles and the volume it generated at thirty degree celsius and 1 atm. Lastly , the amount of the given gas was dissolved in water in the given pressure and temperature.
1
Expert's answer
2021-08-09T07:03:02-0400

Molar Mass of iron = 55.485

1/55.485

= 0.018

Molar Mass of HCl = 36.4

= 3 × 36.4 = 109.2

= 20/109.2

= 0.183 moles

Fe + 10HCl-->> Fe10HCl